OpenStudy (anonymous):

Solve the separable equation: dx/dt - x^3 = x

OpenStudy (amistre64):

since its separable; just separate it ...

OpenStudy (amistre64):

might have to delve into partial fractions for it

OpenStudy (lgbasallote):

and because it's an equation equat it?

OpenStudy (anonymous):

i will write up an equation, give me a minute

OpenStudy (anonymous):

\[\int\limits_{}^{} \frac{ dx }{ dt } = \int\limits_{}^{} x+x^3\] \[\int\limits_{}^{} \frac{ dx }{ x+x^3 } = \int\limits_{}^{} dt\] \[-\frac{ \ln(x^2+1) }{ 2 } + \ln \left| x \right| = t\]

OpenStudy (anonymous):

oops i forgot the + c (constant)

OpenStudy (amistre64):

how did you go from the 2nd to 3rd equation?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=integrate+1%2F%28x^3%2Bx%29 :P lazy write equation out...

OpenStudy (anonymous):

gonna go sleep goodnight people

OpenStudy (amistre64):

pleasant dreams .....

OpenStudy (anonymous):

thanks :D

OpenStudy (anonymous):

\[\frac{1}{x+x^3}=\frac{1}{x(1+x^2)}=\frac{A}{x}+\frac{Bx+C}{1+x^2}\]multiply by \(x\)\[\frac{1}{1+x^2}=A+x\frac{Bx+C}{1+x^2}\]let \(x=0\)\[A=1\]------------------------ multiply by \(1+x^2\)\[\frac{1}{x}=\frac{A(1+x^2)}{x}+Bx+C\]let \(x=i\)\[-i=Bi+C\]\[B=-1 , C=0\]-------------------------\[\frac{1}{x+x^3}=\frac{1}{x}+\frac{-x}{1+x^2}\]

OpenStudy (anonymous):

Thank you so much, SNSDYoona. I'm sorry it took me so long to see this. You helped me out greatly. I appreciate it.